# Locality Sensitive Sparse Encoding for Learning World Models Online

### v2

@@ 1 Introduction
Inspired by Knoll & de Freitas (2012), we propose an expressive sparse non-linear feature representation which we call locality sensitive sparse encoding. Our encoder generates high-dimensional sparse features with random projection and soft binning. Exploiting the sparsity, we further develop an efficient algorithm for online model learning, which only updates a small subset of weights while continually tracking a solution to the FTL objective. World models learned online with our method are resilient to forgetting compared to those based on NNs (see the bottom row of Figure 1 for intuition). We empirically validate the representational advantage of our encoding over other non-linear features, and demonstrate that our method outperforms NNs in online supervised learning (Orabona, 2019; Hoi et al., 2021) as well as model-based reinforcement learning (Sutton, 1990) with models learned online.
@@ 3.2 Feature representation
To give a concrete example, if we set $\rho=3$ and $\lambda=10$, then $\|\phi({\mathbf{x}})\|_{0}$ is at most $8\kappa$ while the dimension of $\phi({\mathbf{x}})$ can be as high as $1000\kappa$. The high-dimensional feature brings us more fitting capacity, and the bounded sparsity allows us to do efficient model update with constant overhead, as we will show in the following section.
@@ 3.3 Efficient sparse incremental update
A no-regret world model after $t$ steps of interaction can be updated online by computing the weights ${\mathbf{W}}_{t+1}$ as in Equation 3, which can be solved in closed form using the normal equation: ${\mathbf{W}}_{t+1}={\mathbf{A}}_{t}^{-1}{\mathbf{B}}_{t}$, where ${\mathbf{A}}_{t}={\mathbf{\Phi}}_{t}^{\top}{\mathbf{\Phi}}_{t}$ and ${\mathbf{B}}_{t}={\mathbf{\Phi}}_{t}^{\top}{\mathbf{Y}}_{t}$ are two memory matrices. Hence, we obtain a simple algorithm (Algorithm 1) to learn the model online at every time step. Algorithm 1 Online model learning 1:$t=0,{\mathbf{A}}_{0}\in\mathbb{R}^{D\times D}=\textbf{0},{\mathbf{B}}_{0}\in\mathbb{R}^{D\times S}=\textbf{0}$ 2:while True do 3:     $t\leftarrow t+1$ 4:     ${\mathbf{A}}_{t}\leftarrow{\mathbf{A}}_{t-1}+{\phi({\mathbf{x}}_{t})}{\phi({\mathbf{x}}_{t})}^{\top}$ 5:     ${\mathbf{B}}_{t}\leftarrow{\mathbf{B}}_{t-1}+{\phi({\mathbf{x}}_{t})}{\mathbf{y}}_{t}^{\top}$ 6:     ${\mathbf{W}}_{t+1}\leftarrow{\mathbf{A}}_{t}^{-1}{\mathbf{B}}_{t}$ 7:end while Algorithm 2 Sparse online model learning 1:$t=0,{\mathbf{A}}_{0}\in\mathbb{R}^{D\times D}=\textbf{0},{\mathbf{B}}_{0}\in\mathbb{R}^{D\times S}=\textbf{0}$ 2:while True do 3:     $t\leftarrow t+1$ 4:     $s\leftarrow\texttt{nonzero\_index}(\phi({\mathbf{x}}_{t}))$ 5:     ${\mathbf{A}}_{t,{\color[rgb]{0.6,0,0.4}ss}}\leftarrow{\mathbf{A}}_{t-1,{\color[rgb]{0.6,0,0.4}ss}}+\phi_{\color[rgb]{0.6,0,0.4}s}({\mathbf{x}}_{t})\phi_{\color[rgb]{0.6,0,0.4}s}({\mathbf{x}}_{t})^{\top}$ 6:     ${\mathbf{B}}_{t,{\color[rgb]{0.6,0,0.4}s}}\leftarrow{\mathbf{B}}_{t-1,{\color[rgb]{0.6,0,0.4}s}}+\phi_{\color[rgb]{0.6,0,0.4}s}({\mathbf{x}}_{t}){\mathbf{y}}_{t}^{\top}$ 7:     ${\mathbf{W}}_{t+1,{\color[rgb]{0.6,0,0.4}s}}\leftarrow{\mathbf{A}}_{t,{\color[rgb]{0.6,0,0.4}ss}}^{-1}({\color[rgb]{0.6,0,0.4}{\mathbf{B}}_{t,s}-{\mathbf{A}}_{t,s{\overline{s}}}{\mathbf{W}}_{t,{\overline{s}}}})$ 8:end while
@@
Updating the memory matrices ${\mathbf{A}}_{t},{\mathbf{B}}_{t}$ is relatively cheap, but computing weights (line $6$ of Algorithm 1) involves matrix inversion, which is computationally expensive. Though ${\mathbf{W}}_{t+1}$ can be recursively updated using Sherman-Morrison formula (Sherman & Morrison, 1950) as done in Least-Squares TD (Bradtke & Barto, 1996), when the feature becomes very high-dimensional to gain expressiveness, it is still inefficient and may no longer be practical.
@@
Fortunately, the incremental updates ${\phi({\mathbf{x}}_{t})}{\phi({\mathbf{x}}_{t})}^{\top}$ and ${\phi({\mathbf{x}}_{t})}{\mathbf{y}}_{t}^{\top}$ are highly sparse by Losse, so that we can efficiently update a small subset of weights. As shown in Algorithm 2, we first record the indices where the softly binned random features fall into (line $4$), which are used to locate those “activated” entries in the current step and update only a small block of the memory matrices (line $5$-$6$). Most importantly, the inversion for computing the weights can be conducted on a much smaller sub-matrix efficiently while ensuring optimality (line $7$). We next derive how our sparse weight update rule solves Equation 3 in a block-wise manner, which yields an optimal fitting for all data including the current one at each update step.
@@
Concluding remark. So far, we have presented an efficient algorithm that learns world models online without forgetting. In our algorithm, linear models are exploited so that we can achieve no-regret online learning with Follow-The-Leader. To enlarge the model capacity, we devise a high-dimensional nonlinear random feature encoding, Losse, turning linear models into universal approximators (Huang et al., 2006). Moreover, the sparsity guarantee of Losse permits efficient updates of the shallow but wide model. We name our method Losse-FTL. We further note that feature sparsity has been utilized to mitigate catastrophic forgetting (McCloskey & Cohen, 1989; French, 1991; Liu et al., 2019; Lan & Mahmood, 2023). Our work differs from them in that not only does sparsity help reduce feature interference, but the incremental closed-form solution also guarantees the optimal fitting of all observed data, thus eliminating forgetting. More discussion on feature sparsity can be found in Appendix A, and the implementation details of Losse-FTL can be found in Appendix D.
@@ 4.1 Comparing feature representations
We test different encoding methods on an image denoising task, where the inputs are MNIST (Deng, 2012) images with Gaussian noise, and the outputs are clean images. The flattened noisy images are first encoded by different methods to produce feature vectors, on which a linear layer is applied to predict the clean images of the same size. We use mini-batch stochastic gradient descent to optimize the weights of the linear layer until convergence. To keep the online update efficiency similar, we keep the same number of non-zero entries for all encoding methods. The mean squared errors on the test set for different patch sizes are reported in Table 1. We observe that Losse achieves the lowest error across all patch sizes, justifying its representational strength over other non-linear feature encoders. When compared with the NN baseline, which is a strong function approximator but not an efficient online learner, linear models with Losse work better when the patch size is $36$ or smaller, suggesting its sufficient capacity for problems with a moderate number of dimensions, such as locomotion or robotics (Todorov et al., 2012; Zhu et al., 2020). See Appendix B.1 for more experimental details.
@@ 4.2 Online learning with covariate shift
As proven in Pan et al. (2021), $X_{t}$ can share the same equilibrium distribution but possess different levels of temporal correlation if $\beta,c$, and $\sigma$ are properly chosen. It turns out the three scalars can be uniquely determined by a single parameter $d\in[0,1)$, which we call correlation level. When $d=0$, the generated data recovers i.i.d. property. Given the data stream $\{(x_{t},y_{t})\}_{t\in\mathbb{N}}$, we fit our model online and measure the mean squared error at the end of learning on a holdout test set, which contains independent samples across all $S_{t}$. As a comparison, we use neural networks to fit online as well as in batch and plot the errors under different correlation levels in Figure 4. The results show that our method outperforms neural networks in two aspects. First, when the data is i.i.d. ($d=0$), Losse-FTL achieves a much lower error than NN-online. This indicates neural networks with gradient descent have low sample efficiency, even on stationary data. Training NNs using batch samples alleviates the issue and reaches a similar performance to ours. Second, our model consistently attains very low error across all correlation levels, showing its capacity of non-forgetting online learning. In contrast, neural networks learned with both batch and online updates incur increasingly high errors when $d$ is large, indicating catastrophic forgetting in the presence of data nonstationarity. More experimental details can be found in Appendix B.2.
@@ 5.1 Settings and baselines
We employ the Dyna MBRL architecture and restrict our study to the model learning part, keeping the base agent’s value (and policy) learning untouched. We compare Losse-FTL world models learned online with their NN-based counterparts, using the same model-free base agent. We introduce the baselines below. The algorithm and more experimental details can be found in Appendix C.
@@ Appendix A On feature sparsity
Appendix B Supervised learning details
In this section, we provide more details regarding the supervised learning experiments on image denoising (Appendix B.1) and piecewise random walk (Appendix B.2).
B.1 Image denoising
@@
B.2 Piecewise random walk
@@
Appendix C Reinforcement learning details
@@
Appendix D Implementation details
@@
We provide more implementation details of Losse-FTL in this section. Referring to Figure 3, the input ${\mathbf{x}}_{t}$ is first pre-processed to be bounded between $[-3,3]$, and then projected with a matrix containing values sampled from an isotropic Gaussian with $\mathbf{\mu}=\mathbf{0}$ and $\mathbf{\Sigma}=\frac{1}{c}\mathbf{I}$, where $c$ is the fan-in dimension, i.e., $S+A$. This ensures the projected values $\sigma({\mathbf{x}}_{t})$ are bounded with high probability and facilitates a more even bin utilization. In computing the sparse update at line 7 of Algorithm 2, we use $({\mathbf{A}}_{t,ss}+\varepsilon{\mathbf{I}})^{-1}$ for some small $\varepsilon>0$ to ensure its invertibility. We provide Jax-based Python codes in Figure 10.
import math
from typing import NamedTuple, Tuple
import chex
import jax
import jax.numpy as jnp
import numpy as np
class LosseParams(NamedTuple):
    count: jax.Array
    projection: jax.Array
    xtx: jax.Array
    xty: jax.Array
    w: jax.Array
def _to_1d_index(indices, offsets, n_feat, bin_dim, n_bins):
    ”””Compute the flattened index into the weight matrix.”””
    n_grids_per_lsh = (n_bins + 1) ** bin_dim
    indices = jnp.reshape(indices, (-1, bin_dim, n_feat))
    offsets = jnp.reshape(offsets, (-1, bin_dim, n_feat))
    indices = jnp.stack([indices, indices + 1], axis=-1)  # [-1, bin_dim, n_feat, 2]
    values = jnp.stack([1.0 - offsets, offsets], axis=-1)  # [-1, bin_dim, n_feat, 2]
    multiplier = jnp.power(n_bins + 1, jnp.arange(bin_dim - 1, -1, -1))
    indices *= multiplier[:, None, None]
    # shape = (-1, n_feat, ) + (2,) * bin_dim
    shape_suffix = [tuple(*p) for p in np.split(np.eye(bin_dim, dtype=np.int32) + 1, bin_dim)]
    indices = sum(jnp.reshape(indices[:, i], (-1, n_feat, *suffix)) for i, suffix in enumerate(shape_suffix))
    values = math.prod(jnp.reshape(values[:, i], (-1, n_feat, *suffix)) for i, suffix in enumerate(shape_suffix))
    # both indices and values has the shape (-1, n_feat, *(2,)*bin_dim) now.
    indices += jnp.expand_dims(
        n_grids_per_lsh * jnp.arange(n_feat), axis=tuple(range(-bin_dim, 1, 1))
    )  # expand 1 dim in the front and bin_dim in the back.
    indices = jnp.reshape(indices, (-1, n_feat * 2**bin_dim))
    values = jnp.reshape(values, (-1, n_feat * 2**bin_dim))
    return indices, values
class Losse:
    ”””Linear regressor with LOcality Sensitive Sparse Encoding (Losse).
    We update the linear weights online sparsely following Algorithm.2 in the paper, i.e., computing the incremental closed-form solution based on newly incoming data points.
    def __init__(
        self,
        inout_dims: Tuple[int, int],
        num_features: int,
        num_bins: int,
        bin_dim: int,
        eps: float,
    ) -> None:
        self.num_features = num_features
        self.num_bins = num_bins
        self.bin_dim = bin_dim
        self.inout_dims = inout_dims
        self.eps = eps
        n_edges = num_bins + 1
        n_grids_per_lsh = (n_edges + 1) ** bin_dim
        self.d = n_grids_per_lsh * num_features
    def init(self, rng: jax.random.PRNGKey) -> LosseParams:
        input_dim = self.inout_dims[0]
        output_dim = self.inout_dims[1]
        std = 1 / jnp.sqrt(input_dim)
        projection = std * jax.random.truncated_normal(
            rng,
            -2,
            2,
            (input_dim, self.num_features * self.bin_dim),
        return LosseParams(
            count=jnp.array(0, dtype=jnp.int64),
            projection=projection,
            xtx=jnp.zeros((self.d * self.d,), dtype=projection.dtype),
            xty=jnp.zeros((self.d, output_dim), dtype=projection.dtype),
            w=jnp.zeros((self.d, output_dim), dtype=projection.dtype),
    def update(
        self,
        params: LosseParams,
        x: jax.Array,
        y: jax.Array,
    ) -> LosseParams:
        chex.assert_tree_shape_prefix((x, y), (1,))  # assert non-batched
        indices, values = self._indices_and_values(params.projection, x)
        params = self._update_memory(params, indices, values, y)
        params = self._update_w(params, indices)
        return params
    def predict(self, params: LosseParams, x: jax.Array):
        indices, values = self._indices_and_values(params.projection, x)
        output = params.w[indices] * values[…, None]
        return output.sum(1)
    def _indices_and_values(
        self,
        projection: jax.Array,
        x: jax.Array,
        h = jnp.matmul(x, projection)
        h = jax.nn.sigmoid(h)
        h = jnp.clip(h, 0, 1) * self.num_bins
        indices = jnp.floor(h).astype(jnp.int32)
        offsets = h - indices
        indices, values = _to_1d_index(
            indices,
            offsets,
            self.num_features,
            self.bin_dim,
            self.num_bins,
        return indices, values
    def _update_memory(
        self,
        params: LosseParams,
        indices: jax.Array,
        values: jax.Array,
        y: jax.Array,
    ) -> LosseParams:
        chex.assert_equal_shape_prefix((indices, values, y), prefix_len=1)
        xtx_indices = (indices * self.d)[:, :, None] + indices[:, None, :]
        xtx_indices = xtx_indices.flatten()
        xty_indices = indices.flatten()
        xtx_updates = values[:, :, None] * values[:, None]
        xtx_updates = xtx_updates.flatten()
        xty_updates = values[:, :, None] * y[:, None, :]
        xty_updates = xty_updates.reshape(-1, y.shape[-1])
        return params._replace(
            xtx=params.xtx.at[xtx_indices].add(xtx_updates),
            xty=params.xty.at[xty_indices].add(xty_updates),
            count=params.count + y.shape[0],
    def _update_w(
        self,
        params: LosseParams,
        indices: jax.Array,
    ) -> LosseParams:
        indices = indices.flatten()
        sub_indices = (indices * self.d)[:, None] + indices[None, :]
        sub_xtx = jnp.reshape(params.xtx[sub_indices], [indices.shape[0]] * 2)
        sub_xty = params.xty[indices]
        a = sub_xtx
        sub_xtxw = jnp.matmul(
            jnp.reshape(params.xtx, (self.d, self.d))[indices],
            params.w,
        b = sub_xty - sub_xtxw + jnp.matmul(sub_xtx, params.w[indices])
        a_norm = a / params.count + self.eps * jnp.eye(len(a))
        b_norm = b / params.count
        solution = jnp.linalg.solve(a_norm, b_norm)
        return params._replace(w=params.w.at[indices].set(solution))
losse = Losse(
    inout_dims=(1, 1),
    num_features=50,
    num_bins=10,
    bin_dim=2,
    eps=1e-5,
losse.init = jax.jit(losse.init)
losse.update = jax.jit(losse.update, donate_argnums=(0,))  # donate to avoid copy
losse.predict = jax.jit(losse.predict)
Jax-based Python codes for learning an online linear regressor with Losse-FTL.

### v4

@@ 1 Introduction
Inspired by Knoll & de Freitas (2012), we propose an expressive sparse non-linear feature representation which we call locality sensitive sparse encoding. Our encoder generates high-dimensional sparse features with random projection and soft binning. Exploiting the sparsity, we further develop an efficient algorithm for online model learning, which only updates a small subset of weights while continually tracking a solution to the FTL objective. World models learned online with our method are resilient to forgetting compared to those based on NNs (see the bottom row of Figure 1 for intuition). We empirically validate the representational advantage of our encoding over other non-linear features, and demonstrate that our method outperforms NNs in the online supervised learning setting (Orabona, 2019; Hoi et al., 2021) as well as the model-based reinforcement learning setting (Sutton, 1990) with models learned online.
@@ 3.2 Feature representation
To give a concrete example, if we set $\rho=3$ and $\lambda=10$, then $\|\phi({\mathbf{x}})\|_{0}$ is at most $8\kappa$ while the dimension of $\phi({\mathbf{x}})$ can be as high as $1000\kappa$. The high-dimensional feature brings us more fitting capacity, while the bounded sparsity permits efficient model updates with constant overheads, as we will show in the following section.
@@ 3.3 Efficient sparse incremental update
A no-regret world model after $t$ steps of interaction can be updated online by computing the weights ${\mathbf{W}}_{t+1}$ (Equation 3), which can be solved in closed form using the normal equation: ${\mathbf{W}}_{t+1}={\mathbf{A}}_{t}^{-1}{\mathbf{B}}_{t}$, where ${\mathbf{A}}_{t}={\mathbf{\Phi}}_{t}^{\top}{\mathbf{\Phi}}_{t}$ and ${\mathbf{B}}_{t}={\mathbf{\Phi}}_{t}^{\top}{\mathbf{Y}}_{t}$ are two memory matrices. Hence, we obtain a simple algorithm (Algorithm 1) to learn the model online at every time step. Algorithm 1 Online model learning 1:$t=0,{\mathbf{A}}_{0}\in\mathbb{R}^{D\times D}=\textbf{0},{\mathbf{B}}_{0}\in\mathbb{R}^{D\times S}=\textbf{0}$ 2:while True do 3:     $t\leftarrow t+1$ 4:     ${\mathbf{A}}_{t}\leftarrow{\mathbf{A}}_{t-1}+{\phi({\mathbf{x}}_{t})}{\phi({\mathbf{x}}_{t})}^{\top}$ 5:     ${\mathbf{B}}_{t}\leftarrow{\mathbf{B}}_{t-1}+{\phi({\mathbf{x}}_{t})}{\mathbf{y}}_{t}^{\top}$ 6:     ${\mathbf{W}}_{t+1}\leftarrow{\mathbf{A}}_{t}^{-1}{\mathbf{B}}_{t}$ 7:end while Algorithm 2 Sparse online model learning 1:$t=0,{\mathbf{A}}_{0}\in\mathbb{R}^{D\times D}=\textbf{0},{\mathbf{B}}_{0}\in\mathbb{R}^{D\times S}=\textbf{0}$ 2:while True do 3:     $t\leftarrow t+1$ 4:     $s\leftarrow\texttt{nonzero\_index}(\phi({\mathbf{x}}_{t}))$ 5:     ${\mathbf{A}}_{t,{\color[rgb]{0.6,0,0.4}ss}}\leftarrow{\mathbf{A}}_{t-1,{\color[rgb]{0.6,0,0.4}ss}}+\phi_{\color[rgb]{0.6,0,0.4}s}({\mathbf{x}}_{t})\phi_{\color[rgb]{0.6,0,0.4}s}({\mathbf{x}}_{t})^{\top}$ 6:     ${\mathbf{B}}_{t,{\color[rgb]{0.6,0,0.4}s}}\leftarrow{\mathbf{B}}_{t-1,{\color[rgb]{0.6,0,0.4}s}}+\phi_{\color[rgb]{0.6,0,0.4}s}({\mathbf{x}}_{t}){\mathbf{y}}_{t}^{\top}$ 7:     ${\mathbf{W}}_{t+1,{\color[rgb]{0.6,0,0.4}s}}\leftarrow{\mathbf{A}}_{t,{\color[rgb]{0.6,0,0.4}ss}}^{-1}({\color[rgb]{0.6,0,0.4}{\mathbf{B}}_{t,s}-{\mathbf{A}}_{t,s{\overline{s}}}{\mathbf{W}}_{t,{\overline{s}}}})$ 8:end while
@@
Updating the memory matrices ${\mathbf{A}}_{t},{\mathbf{B}}_{t}$ is relatively cheap, but computing weights (line $6$ of Algorithm 1) involves matrix inversion, which is computationally expensive. Though ${\mathbf{W}}_{t+1}$ can be recursively updated using Sherman-Morrison formula (Sherman & Morrison, 1950) as done in Least-Squares TD (Bradtke & Barto, 1996), when the feature becomes very high-dimensional to gain expressiveness, it can become inefficient or even impractical.
@@
Fortunately, the incremental updates ${\phi({\mathbf{x}}_{t})}{\phi({\mathbf{x}}_{t})}^{\top}$ and ${\phi({\mathbf{x}}_{t})}{\mathbf{y}}_{t}^{\top}$ are highly sparse by Losse, so that we can efficiently update a small subset of weights. As shown in Algorithm 2, we first record the indices where the softly binned random features fall into (line $4$), which are used to locate those “activated” entries in the current step and update only a small block of the memory matrices (line $5$-$6$). Most importantly, the matrix inversion for updating the weight subset can be conducted on a much smaller sub-matrix efficiently while ensuring optimality (line $7$). The cardinality of the index set $s$ is bounded by the number of non-zero features ($\kappa 2^{\rho}$), giving constant compute cost regardless of the number of data samples. We next derive how our sparse weight update rule solves Equation 3 in a block-wise manner, which yields an optimal fitting for all data including the current one at each update step.
@@
Concluding remark. So far, we have presented an efficient algorithm that learns world models online without forgetting. In our algorithm, linear models are exploited so that we can achieve no-regret online learning with Follow-The-Leader. To enlarge the model capacity, we devise a high-dimensional nonlinear random feature encoding, Losse, turning linear models into universal approximators (Huang et al., 2006). Moreover, the sparsity guarantee of Losse permits efficient updates of the shallow but wide model. We name our method Losse-FTL. We note that feature sparsity has been utilized to mitigate catastrophic forgetting for decades (McCloskey & Cohen, 1989; French, 1991; Liu et al., 2019; Lan & Mahmood, 2023). Our work differs from them in that not only does sparsity help reduce feature interference, but the incremental closed-form solution also guarantees the optimal fitting of all observed data, thus eliminating forgetting. More discussion on feature sparsity can be found in Appendix A. Another related line of research to ours is analytic class-incremental learning (Zhuang et al., 2022; 2023; 2024), which we discuss in Appendix B due to space constraints. The implementation details of Losse-FTL can be found in Appendix E.
@@ 4.1 Comparing feature representations
We test different encoding methods on an image denoising task, where the inputs are MNIST (Deng, 2012) images with Gaussian noise, and the outputs are clean images. The flattened noisy images are first encoded by different methods to produce feature vectors, on which a linear layer is applied to predict the clean images of the same size. We use mini-batch stochastic gradient descent to optimize the weights of the linear layer until convergence. To keep the online update efficiency similar, we keep the same number of non-zero entries for all encoding methods. The mean squared errors on the test set for different patch sizes are reported in Table 1. We observe that Losse achieves the lowest error across all patch sizes, justifying its representational strength over other non-linear feature encoders. When compared with the NN baseline, which is a strong function approximator but not an efficient online learner, linear models with Losse work better when the patch size is $36$ or smaller, suggesting its sufficient capacity for problems with a moderate number of dimensions, such as locomotion or robotics (Todorov et al., 2012; Zhu et al., 2020). See Appendix C.1 for more experimental details.
@@ 4.2 Online learning with covariate shift
As proven in Pan et al. (2021), $X_{t}$ can share the same equilibrium distribution but possess different levels of temporal correlation if $\beta,c$, and $\sigma$ are properly chosen. It turns out the three scalars can be uniquely determined by a single parameter $d\in[0,1)$, which we call correlation level. When $d=0$, the generated data recovers i.i.d. property. Given the data stream $\{(x_{t},y_{t})\}_{t\in\mathbb{N}}$, we fit our model online and measure the mean squared error at the end of learning on a holdout test set, which contains independent samples across all $S_{t}$. As a comparison, we use neural networks to fit online as well as in batch and plot the errors under different correlation levels in Figure 4. The results show that our method outperforms neural networks in two aspects. First, when the data is i.i.d. ($d=0$), Losse-FTL achieves a much lower error than NN-online. This indicates neural networks with gradient descent have low sample efficiency, even on stationary data. Training NNs using batch samples alleviates the issue and reaches a similar performance to ours. Second, our model consistently attains very low error across all correlation levels, showing its capacity of non-forgetting online learning. In contrast, neural networks learned with both batch and online updates incur increasingly high errors when $d$ is large, indicating catastrophic forgetting in the presence of data nonstationarity. More experimental details can be found in Appendix C.2.
@@ 5.1 Settings and baselines
We employ the Dyna MBRL architecture and restrict our study to the model learning part, keeping the base agent’s value (and policy) learning untouched. We compare Losse-FTL world models learned online with their NN-based counterparts, using the same model-free base agent. We introduce the baselines below. The algorithm and more experimental details can be found in Appendix D.
@@ Appendix A On feature sparsity
Appendix B Analytic continual learning
Our work is closely related to a paradigm called analytic continual learning, where the closed-form solution, if obtainable, is employed in training a continual learner. A recent line of research explores this topic focusing on class-incremental learning (Zhuang et al., 2022; 2023; 2024). In particular, Zhuang et al. (2022) employ a pre-trained neural network to extract feature vectors $\bm{X}_{k}^{(\mathrm{fe})}$ after a process called feature expansion, and then learn the weights of a fully connected layer $\hat{\bm{W}}_{\mathrm{FCN}}$ on the frozen features for class-incremental classification. The weights at round $k$ can be solved analytically in a recursive manner utilizing the Woodbury matrix identity:
$\hat{\bm{W}}_{\mathrm{FCN}}^{(k)}=\left[\begin{array}[]{ll}\hat{\bm{W}}_{\mathrm{FCN}}^{(k-1)}-\bm{R}_{k}\bm{X}_{k}^{(\mathrm{fe})\top}\bm{X}_{k}^{(\mathrm{fe})}\hat{\bm{W}}_{\mathrm{FCN}}^{(k-1)}\bm{R}_{k}\bm{X}_{k}^{(\mathrm{fe})\top}\bm{Y}_{k}^{\mathrm{train}}\end{array}\right],$ (10)
where $\bm{Y}_{k}^{\mathrm{train}}$ is the label for class $k$, and $\bm{R}_{k}=\left(\sum_{i=0}^{k}\bm{X}_{i}^{(\mathrm{fe})\top}\bm{X}_{i}^{(\mathrm{fe})}+\gamma\bm{I}\right)^{-1}$ is the regularized feature auto-correlation matrix (similar to the inverse of ${\mathbf{A}}_{t}={\mathbf{\Phi}}_{t}^{\top}{\mathbf{\Phi}}_{t}$ in our formulation), which also has a recursive form:
$\bm{R}_{k}=\bm{R}_{k-1}-\bm{R}_{k-1}\bm{X}_{k}^{(\mathrm{fe})\top}\left(\bm{I}+\underbrace{\bm{X}_{k}^{(\mathrm{fe})}\bm{R}_{k-1}\bm{X}_{k}^{(\mathrm{fe})\top}}_{N_{k}\times N_{k}}\right)^{-1}\bm{X}_{k}^{(\mathrm{fe})}\bm{R}_{k-1}.$ (11)
Though Zhuang et al. (2022) have achieved state-of-the-art class-incremental learning performance, there are a few limitations. First, a deep network must be pre-trained and frozen to serve as an expressive feature encoder, which is simply infeasible in online settings. Second, Equation 11 involves inverting a matrix whose size grows with the number of samples, limiting its scalability. Third, the learning system requires the knowledge of task boundaries for matrix decomposition, which may not be accessible in many real-world scenarios. In contrast, our work develops an expressive feature encoder that does not rely on pre-training, proposes a constant-time update rule exploiting the feature sparsity, and applies to reinforcement learning where the nonstationarity appears without explicit task boundaries.
Appendix C Supervised learning details
In this section, we provide more details regarding the supervised learning experiments on image denoising (Appendix C.1) and piecewise random walk (Appendix C.2).
C.1 Image denoising
@@
C.2 Piecewise random walk
@@
Appendix D Reinforcement learning details
@@
Appendix E Implementation details
@@
We provide more implementation details of Losse-FTL in this section. Referring to Figure 3, the input ${\mathbf{x}}_{t}$ is first pre-processed to be bounded between $[-3,3]$, and then projected with a matrix containing values sampled from an isotropic Gaussian with $\mathbf{\mu}=\mathbf{0}$ and $\mathbf{\Sigma}=\frac{1}{c}\mathbf{I}$, where $c$ is the fan-in dimension, i.e., $S+A$. This ensures the projected values $\sigma({\mathbf{x}}_{t})$ are bounded with high probability and facilitates a more even bin utilization. In computing the sparse update at line 7 of Algorithm 2, we use $({\mathbf{A}}_{t,ss}+\varepsilon{\mathbf{I}})^{-1}$ for some small $\varepsilon>0$ to ensure its invertibility. We provide Jax-based Python codes in Appendix E.
[…]
[frame=single,breaklines=true,numbers=left,fontsize=]python./assets/losse.py {listing} Jax-based Python codes for learning an online linear regressor with Losse-FTL.